OpenStudy (anonymous):

evaluate the integral to 3 decimal places

OpenStudy (anonymous):

\[\int\limits_{-3}^{3}\sqrt{9-x ^{2}}\]

OpenStudy (anonymous):

integral of the form sqrt(a^2 - x^2) = (x/2)*sqrt(a^2 - x^2) + ((a^2)/2)*arcsin(u/a) + constant, evaluated at from -3 to 3 = (9*pi)/2 to 3 decimal places = 14.137

OpenStudy (anonymous):

can u use \[\int\limits_{}^{}(ax+b)^{n}dx= (ax+b)^{n+1}/a(n+1) +c\]

OpenStudy (anonymous):

i haven't used integration with trigonometric identities yet

OpenStudy (anonymous):

the problem is that i get zero when i integrate using that method

OpenStudy (anonymous):

and i know its not possible to get zero area in that problem lol

OpenStudy (anonymous):

ax +b will not work because you have the form b - ax^2.

OpenStudy (anonymous):

oh i see do you no what form i should use exlcuding the use of trig forms

OpenStudy (anonymous):

i didnt learn integration using trigonometric identities

OpenStudy (dumbcow):

i would use trig substitution x=3sin(u)

OpenStudy (dumbcow):

im not sure how else to do it

OpenStudy (anonymous):

so there isnt another method?

OpenStudy (anonymous):

oh i see

OpenStudy (dumbcow):

well because its a definite integral you could approximate area using graphing software or sum of rectangles

OpenStudy (anonymous):

it asks to evaluate the integral does that mean i cant use the area formulae of a circle

OpenStudy (dumbcow):

hmm i didnt notice that before yeah this is just the area of the upper part of a circle of radius 3

OpenStudy (dumbcow):

so A= pi*r^2/2

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

if it asks to evaluate it doesn't necessarily mean through the use of integration does it

OpenStudy (dumbcow):

it can but in this case there is a more practical and simple method

OpenStudy (anonymous):

i see well thankyou for your help :)

OpenStudy (dumbcow):

\[\int\limits \sqrt{a^2 - x^2} dx = \frac{1}{2} x \sqrt{a^2-x^2} +\frac{1}{2}a^2\tan^{-1}\frac{x}{\sqrt{a^2-x^2}}+C\] http://integral-table.com/

OpenStudy (anonymous):

so that is the form used to integrate that question. I havent learnt it yet so i will assume that using area formulae and a graphy is the way to go lol:)

OpenStudy (dumbcow):

i agree